• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 08:04:47 on June22,2000

    Entry number 44647

    keyword=Shift Summary



    Until 1:00, finish data taking at (-90,180). Final accumulated charge at this kinematics is 10.1907 C.
    1:13 Start quick test run to check E-arm VDC efficiencies at 5uA. Run# 2424.
    1:16 Dto, but lower prescale factors. Run# 2425.
    1:22 Setting up for pointing runs.
    1:27 Pointing run w/ 10cm dummy, run# 2426
    1:35 15cm dummy, run# 2427
    1:41 4cm dummy, run# 2428
    1:50 carbon target, run# 2429

    some time while hrsh is being changed by other losers on shift:
    HARP scan #943, coda run 2430

    2:05 End harp scan. Setting up for new kinematics (50,180).
    2:35 New kinematics set. Run 2431 is junk.
    2:41 Carbon pointing run, run# 2432
    2:49 4cm dummy, run# 2433
    2:57 15cm dummy, run# 2434
    3:07 10cm dummy, run# 2435
    3:12 Moved FPP 1.5, 3, 9in carbon doors in, as per run plan
    Setting up for production running.
    3:18 Run 2436 is raster check. Prescale factors are all 1 except PS5=65000.
    3:20 Asking for 70uA.
    3:27 Trigger efficiency test, 71.5uA, run# 2437.
    3:28 Beam died for some 30 seconds during this run.
    3:31 End trigger check
    3:33 Start production for (50,180) kinematics, run# 2438
    4:47 Opened collimators on both spectrometers. Next run# 2440
    7:40 Jim Kelly calls, alerts us to a problem with the BCMs (see halog 44646)

    Production:
    Run number Events Comments
    2438 1.3M 6msr collimators
    2439 1.3M
    2440 1.3M Collimators now "open"
    2441 750k
    2442 1.25M T2/3 prescale factors changed. Ended by CODA crash.
    2443 -0- junk
    2444 1.3M
    2445 1.3M